# Frame of Reference: Addressing the Challenges of Common Ground Representation in Situational Dialogs

### v1

@@ Abstract
Common ground plays a critical role in situated spoken dialogs, where interlocutors must establish and maintain shared references to entities, events, and relations to sustain coherent interaction in a shared space and over time. With the increasing presence of embodied conversational agents and social robots, the ability to correctly ground this kind of conversational content in order to refer back later also becomes important for dialog systems. Prior studies have demonstrated that LLMs are capable of performing certain grounding acts like acknowledgments. However, relatively little work has investigated their capacity to leverage the grounded information, like in complex scenarios involving space and time (e.g., "let’s go to that café near the park we went to yesterday"). To that end, in this work, we evaluate a model’s ability to establish common ground by utilizing these "relational references" in the dynamic and shared environments of situated dialogs. We then test multiple methods for representing common ground and further propose approaches to improve their performance.
@@ 4.2 Resource Constrained Setting
We compare three commonly used methods for representing common ground under these constraints: summarization, utterance chunking, and ontology creation.
Summarization: A method shown by Padmakumar et al. (2023) to be a highly effective representation of open-dialog context for transformer-based models. Its utility is further underscored by its adoption in modern LLM chatbots for personalization Zhang et al. (2024) by summarizing each session separately. In this scheme, $W$ creates a summary ($s_{t}$) of the whole dialog, which is stored, retrieved by $R$, and used by $G$ to form the next utterance.
@@ 4.3 Evaluation
Table 3 details the performance of various resource-constrained representation techniques. Due to space constraints, we primarily report results for the Llama3.1-8B model; results for the Gemma2 2B and 9B models are provided in Appendix F. The general trend across all models remained consistent. We did not test larger models due to resource limitations. All the representation techniques yielded poorer performance compared to the baseline, a likely consequence of information loss during the extraction of relevant information. We noticed that sparse embedding (BM25) demonstrated a slight advantage over dense embedding, which may be attributed to its enhanced capability in named-entity retrieval. The agentic ontology-based system outperformed other resource-constrained representation techniques. While it performed relatively well on the Meetup dataset, it struggled with the STD dataset, often merging the representations of images of both participants due to their similarity. However, its better performance suggests that a multi-step retrieval process, particularly one that explicitly extracts speaker-specific information regarding entities and events, helps the LLM understand the context better.
@@ 5.1 Synthetic Data
Given that off-the-shelf LLMs lack proficiency in spoken dialog, we designed a systematic, multi-phase synthetic data generation framework with an "environment-first, dialog-later" approach. Inspired by procedural content generation Smith (2015) and recent advances in synthetic dialog generation Suresh et al. (2025), our framework is engineered to produce a grounded situated dialog corpus. The process is structured into three phases as shown in Figure 2:
@@
To assess whether the enhanced contextual understanding of GRPO-trained models improves their representational capabilities, we applied the agentic ontology-based representation using the same GRPO-trained model trained on synthetic Q/A pairs, given its promising results in earlier experiments. As shown in Table 5, the GRPO-trained model outperformed its non-GRPO counterpart on the Meetup dataset but continued to struggle with the Spot-the-Difference dataset. The main issue concerned representing the two images as distinct within the common ground, as the models frequently merged them due to their high degree of similarity. While more carefully designed task-specific prompts could improve performance, these would limit generalizability. Thus, our findings fall short of our broader goal: developing a universal model for a persistent and useful common ground representation capable of resolving complex relational references to it while relying only on spoken utterances and minimal contextual descriptions.
@@ Appendix B Benchmark Creation
Illustrative examples for each test case category are provided in Figures 3, 4, 5, and 6. Specifically, we present two examples (Spatial and Temporal) from the Spot the Difference dialogs and two examples (Attributive and Inferred) from the Meetup dialogs. In these figures, we have highlighted the essential segment of the dialog containing the answer (in purple) to the question (in violet). We also provide the keywords needed for the exact match analysis.
@@ Appendix F Gemma Results
Tables 6 and 7 contain all the results for Gemma2 2B and 9B models. As seen from the Tables, the overall pattern remains the same as compared to Llama 3.1-8B. The 2 Billion model performs relatively poor compared to the 9 Billion model across all the categories. Chunking in general provides the worst results among all the representation techniques suggesting the importance of good abstraction of information that preserves the links between the information across different segments of the conversation. Surprisingly, for the gemma2 models that struggled with longer context of STD, the agentic system performs better than when provided with the full dialog for Meetup test cases. This suggests that providing a structure might help it reason better for longer contexts.
@@ Appendix G Prompts
During testing, we utilized the system prompts shown in Figures 9, 10 and 11. Our approach leverages chain-of-thought prompting (Figure 10), which we found to be more effective than direct questioning. This prompt outlines the game’s rules to help the LLM understand the context. However, it differs from the in-context prompt (Figure 11) by not specifying how to formulate the answer. We have omitted the prompts for our Agentic method from this paper due to the numerous large prompts used in its different phases. Furthermore, we deliberately excluded textual descriptions of the images from the prompts. This ensures that the questions can only be answered using the conversational context, thereby testing the model’s ability to form an accurate common ground representation based on what has been discussed.

### v2

@@ Abstract
Common ground plays a critical role in situated spoken dialogs, where interlocutors must establish and maintain shared references to entities, events, and relations to sustain coherent interaction in a shared space and over time. With the increasing presence of embodied conversational agents and social robots, the ability to correctly ground this kind of conversational content in order to refer back later also becomes important for dialog systems. Prior studies have demonstrated that LLMs are capable of performing certain grounding acts like acknowledgments. However, relatively little work has investigated their capacity to leverage the grounded information, like in complex scenarios involving space and time (e.g., "let’s go to that café near the park we went to yesterday"). To that end, in this work, we evaluate a model’s ability to establish common ground by utilizing these "relational references" in the dynamic and shared environments of situated dialogs. We then test multiple methods for representing common ground and further propose approaches to improve their performance by using reinforcement learning on our synthetically generated dialog data .
@@ 4.2 Resource Constrained Setting
We compare three commonly used common ground representation methods under these constraints.
Summarization: A method shown by Padmakumar et al. (2023) to be an effective representation of open-dialog context for transformer-based models. Its utility is further underscored by its adoption in modern LLM chatbots for personalization Zhang et al. (2024) by summarizing each session separately. In this scheme, $W$ creates a summary ($s_{t}$) of the whole dialog, which is stored, retrieved by $R$, and used by $G$ to form the next utterance.
@@ 4.3 Evaluation
Table 3 details the performance of various resource-constrained representation techniques. Due to space constraints, we primarily report results for the Llama3.1-8B model; results for the Gemma2 2B and 9B models are provided in Appendix F. The general trend across all models remained consistent. We did not test larger models due to resource limitations. All the representation techniques yielded poorer performance compared to the baseline, a likely consequence of information loss during the extraction of relevant information. We noticed that sparse embedding (BM25) demonstrated a slight advantage over dense embedding, which may be attributed to its enhanced capability in named-entity retrieval. The agentic ontology-based system outperformed other resource-constrained representation techniques. While it performed relatively well on the Meetup dataset, it struggled with the STD dataset, often merging the representations of images of both participants due to their similarity. However, its better performance suggests that a multi-step retrieval process, particularly one that explicitly extracts speaker-specific information regarding entities and events, helps the LLM understand the context better. Code link for evaluation.
@@ 5.1 Synthetic Data
Given that off-the-shelf LLMs lack proficiency in spoken dialog, we designed a systematic, multi-phase synthetic data generation framework with an "environment-first, dialog-later" approach. Inspired by procedural content generation Smith (2015) and recent advances in synthetic dialog generation Mohapatra et al. (2021); Suresh et al. (2025), our framework is engineered to produce a grounded situated dialog corpus. The process is structured into three phases as shown in Figure 2:
@@
To assess whether the enhanced contextual understanding of GRPO-trained models improves their representational capabilities, we applied the agentic ontology-based representation using the same GRPO-trained model trained on synthetic Q/A pairs, given its promising results in earlier experiments. As shown in Table 5, the GRPO-trained model outperformed its non-GRPO counterpart on the Meetup dataset but continued to struggle with the Spot-the-Difference dataset. The main issue concerned representing the two images as distinct within the common ground, as the models frequently merged them due to their high degree of similarity. While more carefully designed task-specific prompts could improve performance, these would limit generalizability. These findings therefore suggest that better multi-step reasoning alone does not guarantee a robust, reusable representation of persistent common ground. Here is the link for the synthetic data generation code.
Figure 3: Visual representation of RL for optimizing the externalized common ground representation.
Setting Meetup Spot the Difference
Off-the-shelf Llama [Table 3] 0.40 / 0.36 0.38 / 0.34 0.38 / 0.30 0.24 / 0.22 0.20 / 0.14 0.12 / 0.14 0.20 / 0.24 0.18 / 0.14
Template Constrained 0.30 / 0.22 0.28 / 0.26 0.18 / 0.20 0.20 / 0.16 0.12 / 0.10 0.10 / 0.10 0.18 / 0.16 0.14 / 0.12
Open Ended 0.04 / 0.04 0.02 / 0.04 0.06 / 0.06 0.02 / 0.02 0.02 / 0.04 0.04 / 0.04 0.08 / 0.06 0.00 / 0.00
Table 6: FEM ($\uparrow$) / LLM-as-Judge ($\uparrow$) accuracy for agentic ontology-based representation where the common ground generator is directly trained using RL.
5.2 RL for optimizing the externalized common ground representation
We further investigated whether RL could directly optimize the externalized common ground representation produced by our agentic framework. In this setting, the common ground generator (Writer) serves as the policy: given a dialog, it produces a common ground representation that is then consumed by the Reader and Generator modules to answer an IndiRef question. The correctness of the final answer, as determined by our evaluation pipeline, was used as the reward signal for the representation step (see Figure 3). Training was conducted on the synthetic dialogs introduced above.
We evaluated two variants. The template constrained setting required the model to populate our ontology-style schema. This meant that the overall structure of the representation was fixed in advance, and reinforcement learning could only optimize what information to include within that structure, such as which entities, attributes, relations, and temporal events were most useful for later retrieval. In effect, this setup tests whether RL can make a predefined representation more functionally informative. The open-ended setting removed the schema entirely and asked the model to generate any representation that would allow another model to answer future questions from the shared context. Here, the model was not given an explicit ontology or field structure, and therefore had to decide both what information to preserve and how to organize it. This setup tests whether the model can discover, through reward alone, a representation strategy better suited for downstream retrieval than a hand-designed template.
Contrary to our expectations, both RL settings underperformed the off-the-shelf ontology pipeline (Table 6). The template-constrained variant did not improve over the original system, and the open-ended variant collapsed to near-zero accuracy. Our error analysis suggests that the main bottleneck is the credit assignment problem in multi-step agentic pipelines. Since reward is assigned only from the correctness of the final answer, it does not cleanly reflect the quality of the generated common ground. As a result, training was corrupted by Spurious Rewards, in which poor representations were reinforced because downstream modules still produced the correct answer, and Unjustified Penalties, in which high-quality representations were discouraged because later stages failed to use them correctly. In the open-ended setting, the absence of structural constraints led the model to produce generic summaries rather than the fine-grained relational content needed for IndiRef.
These findings indicate that optimizing externalized common ground via end-to-end RL is substantially harder than improving direct answer generation. They also suggest that for grounding representation, a structured template provides an important inductive bias that cannot be reliably replaced by open-ended generation.
@@ 6 Conclusion and Future Work
Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation [AD011016333] made by GENCI.
@@ Appendix B Benchmark Creation
Illustrative examples for each test case category are provided in Figures 4, 5, 6, and 7. Specifically, we present two examples (Spatial and Temporal) from the Spot the Difference dialogs and two examples (Attributive and Inferred) from the Meetup dialogs. In these figures, we have highlighted the essential segment of the dialog containing the answer (in purple) to the question (in violet). We also provide the keywords needed for the exact match analysis.
@@ Appendix F Gemma Results
Tables 7 and 8 contain all the results for Gemma2 2B and 9B models. As seen from the Tables, the overall pattern remains the same as compared to Llama 3.1-8B. The 2 Billion model performs relatively poor compared to the 9 Billion model across all the categories. Chunking in general provides the worst results among all the representation techniques suggesting the importance of good abstraction of information that preserves the links between the information across different segments of the conversation. Surprisingly, for the gemma2 models that struggled with longer context of STD, the agentic system performs better than when provided with the full dialog for Meetup test cases. This suggests that providing a structure might help it reason better for longer contexts.
@@ Appendix G Prompts
During testing, we utilized the system prompts shown in Figures 10, 11 and 12. Our approach leverages chain-of-thought prompting (Figure 11), which we found to be more effective than direct questioning. This prompt outlines the game’s rules to help the LLM understand the context. However, it differs from the in-context prompt (Figure 12) by not specifying how to formulate the answer. We have omitted the prompts for our Agentic method from this paper due to the numerous large prompts used in its different phases. Furthermore, we deliberately excluded textual descriptions of the images from the prompts. This ensures that the questions can only be answered using the conversational context, thereby testing the model’s ability to form an accurate common ground representation based on what has been discussed.